OpenStudy (anonymous):

Use the order of operations to simplify the expression. Really needing this explained.

OpenStudy (anonymous):

Sure. Everything between ( ) must be worked don first.

OpenStudy (anonymous):

Then, any exponents must be worked on next. Followed by multiplication and division. Followed by addition and subtraction.

OpenStudy (anonymous):

\[(\frac{ 7 }{ 8 }-\frac{ 1 }{ 4 }) \div \frac{ 10 }{ 9 }\]

OpenStudy (anonymous):

The multiplication and division are interchangeable... same for addition and subtraction. For example... 1*2/2 = 2*1/2

OpenStudy (anonymous):

Okay... what would you start with first?

OpenStudy (anonymous):

The parenthesis

OpenStudy (anonymous):

Good! But, in the parenthesis, what operation would you work on first?